# Unsupervised motion segmentation in one go: Smooth long-term model over a video

### v2

@@ 3.2 Loss function
We consider a volume of optical flow fields $f\in\mathbb{R}^{2\times T\times W\times H}$, the spatial grid $\Omega\in\mathbb{R}^{W\times H}$ and $T$ temporal steps. We denote $f(i,t)\in\mathbb{R}^{2}$ the flow associated to site (or point) $i\in\Omega$ ($i=(x,y)$) at time $t$. We assume that we can decompose the flow as a set of $K$ segments, each one exhibiting a coherent motion. Flow vectors within a given segment $k$ are represented by a smooth parametric motion model parametrized by $\vartheta_{k}$. Variable $z_{i,t}$ conveys the motion segmentation, $z_{i,t}^{k}=1$ if site $(i,t)$ belongs to segment $k$. $z$ and $\vartheta=\{\vartheta_{k},k=1\cdots K\}$ are latent variables, and $f$ is the observed data. Following [1], we introduce an approximate distribution over segmentation and motion model parameters:
@@
$\log(\mathcal{D})=\sum_{f\in D}\log p(f)=$ $\sum_{f\in D}(\mathbb{E}_{q(z,\vartheta|f)}[\log\frac{p(f,z,\vartheta)}{q(z,\vartheta|f)}]+KL(q(z,\vartheta|f)||p(z,\vartheta|f))])$
$\geq$ $\sum_{f\in D}(\mathbb{E}_{q(z,\vartheta|f)}[\log p(f|z,\vartheta)]-KL(q(z,\vartheta|f)||p(z,\vartheta))),$ (4)
@@
where we get the Evidence Lower Bound (ELBO). Considering that the flow field can be represented by segment-wise parametric motion models, and assuming conditional independence, we have:
$p(f|z,\vartheta)=\prod_{t=1}^{T}\prod_{k=1}^{K}p(f(t)|\vartheta_{k},z)\propto\prod_{t=1}^{T}\prod_{k=1}^{K}\prod_{i\in\Omega}\exp(-||f(i,t)-\tilde{f}_{S_{n}(\vartheta_{k})}(i,t)||_{1}/\xi_{f,t})^{z_{i,t}^{k}},$ (5)
@@
where $\tilde{f}_{S_{n}(\vartheta_{k})}(i,t)$ is the flow vector given by the parametric motion model of parameters $\vartheta_{k}$ at point $(i,t)$ as defined in Section 3.1 through the spline approximation and $\xi_{f,t}=\sum_{i}||f(i,t)||_{1}$ is a normalizing factor. From this likelihood, we can write the left term of the ELBO as:
$\mathcal{L}_{r}=-\mathbb{E}_{q(z,\vartheta|f)}[\log p(f|z,\vartheta)]$ $=-\sum_{t=1}^{T}\sum_{k}\mathbb{E}_{q}[\log p(f(t)|\vartheta_{k},z)]$
$=\sum_{t=1}^{T}\sum_{i\in\Omega}\sum_{k}g_{\phi}(f)_{i,t}^{k}||f(i,t)-\tilde{f}_{S_{n}(\theta_{k})}(i,t)||_{1}/\xi_{f,t}$ (6)
The term $KL(q(z,\vartheta|f)||p(z,\vartheta)))$ allows us to define a prior over the segmentation. We want to enforce the property that the segmentation labels are temporally consistent, i.e., similar between neighbouring frames. We approximate this term using the temporal consistency loss defined in [4], as it showed good experimental results and was efficient to compute at training time. In addition, we provide a mathematical development that motivates this expression in the Appendix. We have:
$\mathcal{L}_{c}=\frac{1}{2K|\mathcal{I}|}\sum_{i\in\Omega}\sum_{t=2}^{T}\mathbb{1}\Big{[}||f_{i,t}-f_{i,t-1}||_{1}<\lambda\Big{]}\sum_{k=1}^{K}\;|g_{\phi}(f)_{i,t}^{k}-g_{\phi}(f)_{i,t-1}^{k}|,$ (7)
The threshold $\lambda$ allows us to discard occlusion areas. The loss function is thus defined by:
$\sum_{f\in\mathcal{D}}\mathcal{L}(f,\phi,\theta)=\sum_{f\in\mathcal{D}}\mathcal{L}_{r}(f,\phi,\theta)+\beta\mathcal{L}_{c}(f,\phi).$ (8)
@@
We set $\beta=1$. The training procedure alternatively updates $\theta$ and $\phi$ with $\alpha$ as learning rate:
for $\text{for }f\in D:\;\;\theta^{*}=\arg\min_{\theta}L(f,\phi,\theta)(f)\;;\;\phi_{t+1}=\phi_{t}-\alpha\nabla_{\phi}L(f,\phi,\theta^{*})$ (9)
@@ 3.3 Network architecture
The overall architecture of our unsupervised multiple motion segmentation framework is illustrated in Fig.1. It includes two main modules. The first one, taking the flow volume as input, is a 3D U-net [30]. The latent code augmented with embedding positions is directed to the transformer decoder. Then, by cross-attention, the output formed by the volume of segmentation masks is produced. The training of the overall architecture is based on the minimization of the loss function defined below, while the motion model parameters of the segments are given by the B-spline estimation. Temporal consistency is provided by the loss function and the space-time motion models.
@@ 4.1 Implementation details
Following [6, 16, 21, 4, 36], we adopt the RAFT method [32] to compute the optical flow fields. More specifically, we use the RAFT version trained on the MPI Sintel dataset [2]. We downsample the computed flow fields to feed the network with 128$\times$224 vector fields. The output segmentation maps are upsampled to the initial frame size for evaluation. Thus, we can perform efficient training and inference stages. We typically take flow volumes of temporal length $T=9$ at training time. However, at test time, we can process flow volumes of larger temporal length, and even the flow volume of the whole sequence. To compute of the spatio-temporal parametric motion models, $x$ and $y$ coordinates are normalized within $[-1,1]$, and a similar normalization is applied to the $t$ coordinate. The full 12-parameter quadratic motion model is chosen in all the experiments, and we take $\nu=3$ for the frequency factor in the B-spline approximation. We select for each point $(x,y)$ the segment $\hat{k}$ with the highest prediction. In eq.(7), we set $\lambda$ as the $99^{th}$ quantile of the flow differences over the flow volume.
Our LT-MS method is very efficient at test time. The computational time amounts on average to 210 fps on a GPU P100, that is, twice as fast as the ST-MS method [4]. It is certainly due to the long flow sequence given as input to the network, which allows for parallelisation of some heavy computations. In addition, our LT-MS architecture remains lightweight, since it combines only three U-net layers and a transformer decoder on the downsampled feature space. Let us also stress that our LT-MS method does not involve any post-processing at all.
@@ 4.2 Data augmentation and network training
We applied two types of data augmentation dedicated to optical flow. First, we add a global flow to the input flow similarily as in the EM [21] and ST-MS [4] methods. However, in our case, the global flow is given by a full spline-based spatio-temporal motion model whose parameters are chosen at random. The same global flow is added to the flow fields of a given input volume. It allows us to mimic diverse camera motions, enforcing that the motion segments are independent of it. In addition, we corrupt a few flows out of the nine input ones. Thus, we simulate poorly estimated flow fields at some time instants, and the temporal consistency should overcome it.
Our motion segmentation method is fully unsupervised. We never use any manual segmentation annotation. We train our model using the FlyingThings3D (FT3D) dataset [20]. Our (demanding) objective is to demonstrate the high generalization power of our model. We train our model once for all on the optical flow fields computed on FT3D, and then, apply it on new unseen datasets. For hyperparameter setting, we select the stopping epoch from the loss function evaluated on the DAVIS2016 training set. Additional information on the optimization stage in eq.(9) is given in the Appendix. We trained on a single Quadro RTX 6000 for 77h with a batch size of 5 on 32k flow fields computed on FT3D.
@@ 5 Experimental Results
We have carried out experiments on four datasets: DAVIS2016111https://davischallenge.org/index.html [26], SegTrackV2222https://paperswithcode.com/dataset/segtrack-v2-1 [15], FBMS59 [24], and DAVIS2017-motion [5]. Information on the datasets is provided in the Appendix.
@@ 5.1 Ablation study
We have conducted an ablation study to assess three main components of our method LT-MS with four masks ($K=4$), in particular related to the temporal dimension of the problem. We have changed one component at a time as specified hereafter. We use the polynomial space-time quadratic motion model of ST-MS [4] instead of the space-time motion model based on B-splines over the input sequence. We omit the consistency term $\mathcal{L}_{c}$ in the loss function. We just take the convnet without the transformer decoder. All the ablation experiments were run on the three datasets, DAVIS2016, FBMS59, SegTrackV2. Results are collected in Table 1. In addition, we performed them for two input sequence configurations, respectively input sequences of ten flows, and input sequence of 120 flows (in practice, the whole video for the DAVIS2016 dataset). We can observe that the three ablations have almost the same impact on the performance. The three corresponding model components, i.e., the spline-based motion model, the temporal-consistency loss term, and the transformer decoder are thus all beneficial in similar proportions. They are able to handle the temporal dimension of the problem and the temporal motion evolution along the sequence in a compelling way. Admittedly, the contributions of these three components are more significant for the FBMS59 and the SegTrackV2 datasets. However, the dynamic content of the majority of the DAVIS2016 videos, and then, the overall performance score cannot allow to fully appreciate the contributions of the three model components. Yet, they can be acknowledged by visualizing results obtained on some videos of DAVIS2016, as shown in the Appendix.
Table 1: Ablation study for three main components of our method LT-MS ($K=4$) on DAVIS2016, FBMS59 and SegTrackV2. Only one model component is modified at a time. The performance scores are given by the Jaccard index $\mathcal{J}$. We report ablation results with two input-flow sequence length (or cut size), respectively, by dividing the video into pieces of ten successive frames or by considering 120 successive frames (in practice, the whole video for DAVIS2016).
Ablation / Dataset DAVIS2016 FBMS59 SegTrackv2
Cut Size 10 120 10 120 10 120
Full Model LT-MS-K4 74.8 72.4 61.0 58.2 61.3 60.4
Unet3D only 73.0 71.3 56.6 55.5 58.2 57.3
No consistency term $\mathcal{L}_{c}$ 73.5 71.0 57.5 55.5 58.0 57.5
Polynomial space-time quadratic model 73.4 69.8 57.4 54.5 57.8 56.6
5.2 Quantitative and comparative evaluation
We report in Table 2 the results obtained by two versions of our LT-MS method on the three datasets DAVIS2016, SegTrackV2, and FBMS59. LT-MS-K2 performs segmentation with only two masks ($K=2$) while LT-MS-K4 involves four masks ($K=4$). Table 2 collects results obtained by our LT-MS method and other existing unsupervised methods, when available. We follow the categorization proposed in [21] regarding input and training. However, we have added a category w.r.t. the network input for four recent methods, [6, 8, 16, 17], that only use RGB images as imput at test time, the optical flow being only involved in the loss function. Evaluation is performed on the binary ground truth (foreground moving object vs background) for the three datasets. In the Appendix, we explain how we select the segments for the binary evaluation from the multiple motion segments delivered by our LT-MS method. We still put the OCLR method [5] in the category of unsupervised methods, whereas the authors of the DivA method [2] did not. Indeed, OCLR is not fully unsupervised, since it relies on human-annotated sprites to include realistic shapes in the computer-generated data used at training time. We consider the OCLR version taking only optical flow as input. The post-processing added to the CIS method [37], based on Conditional Random Fields (CRF), is an heavy one, which leads most authors to retain only the version without post-processing for a fair comparison.
@@ 5.2 Quantitative and comparative evaluation
Method Training Input DAVIS2016 SegTrack V2 FBMS59
$\mathcal{J}$uparrow $\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{J}$uparrow
Ours LT-MS-K4 Unsupervised Flow 74.8 72.2 61.3 61.0
Ours LT-MS-K2 70.3 68.5 58.6 55.3
ST-MS (4 masks) [4] 73.2 70.3 55.0 59.4
EM [21] 69.3 70.7 55.5 57.8
MoSeg [36] 68.3 66.1 58.6 53.1
FTS [25] 55.8 - 47.8 47.7
[…]
OCLR* [5] (flow only) 72.1 - 67.6 65.4
GWM [6] RGB (Flow in loss) 79.5 - 78.3 77.4
RCF [16] 80.9 - 76.7 69.9
AMD [17] 57.8 - 57.0 47.5
MOD [8] 73.9 - 62.2 61.3
DivA(4$\dagger$ [2] RGB & Flow 72.4 - 64.6 -
[…]
CIS - No Post [37] 59.2 - 45.6 36.8
CIS - With Post [37] 71.5 - 62.0 63.6

### v3

@@ 3.2 Loss function
We consider a volume of optical flow fields $f\in\mathbb{R}^{2\times T\times W\times H}$ on the image grid $\Omega\in\mathbb{R}^{W\times H}$ of width $W$ and height $H$, and over $T$ temporal steps (or, in other words, a sequence of length $T$). We denote $f(i,t)\in\mathbb{R}^{2}$ the flow associated to site (or point) $i\in\Omega$ ($i=(x,y)$) at time $t$. We assume that we can decompose the flow as a set of $K$ segments, each one exhibiting a coherent motion. Flow vectors within a given segment $k$ are represented by a smooth parametric motion model parametrized by $\vartheta_{k}$. Variable $z_{i,t}$ conveys the motion segmentation, $z_{i,t}^{k}=1$ if site $(i,t)$ belongs to segment $k$. $z=\{z_{i,t}^{k},i\in\Omega,t=1,\cdots T,k=1\cdots K\}$ and $\vartheta=\{\vartheta_{k},k=1\cdots K\}$ are the latent variables, and $f$ is the observed data. Following [17], we introduce an approximate distribution over segmentation and motion model parameters:
@@
$\log(\mathcal{D})=$ $\sum_{f\in D}\log p(f)$
$=$ $\sum_{f\in D}\mathbb{E}_{q(z,\vartheta|f)}[\log\frac{p(f,z,\vartheta)}{q(z,\vartheta|f)}]+KL(q(z,\vartheta|f)||p(z,\vartheta|f))$
$\geq$ $\sum_{f\in D}\mathbb{E}_{q(z,\vartheta|f)}[\log p(f|z,\vartheta)]-KL(q(z,\vartheta|f)||p(z,\vartheta)),$ (4)
where we get the Evidence Lower Bound (ELBO) [17].
@@
Considering that the flow field can be represented by segment-wise parametric motion models, and assuming conditional independence, we have:
$p(f|z,\vartheta)$ $=\prod_{t=1}^{T}\prod_{k=1}^{K}p(f(t)|\vartheta_{k},z)$
$\propto\prod_{t=1}^{T}\prod_{k=1}^{K}\prod_{i\in\Omega}\exp(-||f(i,t)-\tilde{f}_{S_{n}(\vartheta_{k})}(i,t)||_{1}/\xi_{f,t})^{z_{i,t}^{k}},$ (5)
@@
where $\tilde{f}_{S_{n}(\vartheta_{k})}(i,t)$ is the flow vector given by the parametric motion model of parameters $\vartheta_{k}$ at point $(i,t)$ as defined in Section 3.1 through the spline approximation and $\xi_{f,t}=\sum_{i\in\Omega}||f(i,t)||_{1}$ is a normalizing factor. From this likelihood, we can write the left term of the ELBO as:
$\mathcal{L}_{r}(f)$ $=-\mathbb{E}_{q(z,\vartheta|f)}[\log p(f|z,\vartheta)]=-\sum_{t,k}\mathbb{E}_{q}[\log p(f(t)|\vartheta_{k},z)]$
$=\sum_{t=1}^{T}\sum_{i\in\Omega}\sum_{k=1}^{K}g_{\phi}(f)_{i,t}^{k}||f(i,t)-\tilde{f}_{S_{n}(\theta_{k})}(i,t)||_{1}/\xi_{f,t}$ (6)
The term $KL(q(z,\vartheta|f)||p(z,\vartheta)))$ in eq.(4) allows us to define a prior over the segmentation. We want to enforce the property that the segmentation labels are temporally consistent, i.e., similar between neighbouring frames. More formally, we express a prior on $z$ as follows, assuming point independence:
$\forall i\in\Omega,\forall t\in\{1\cdots T\},z_{i,t}\in\{1\cdots K\}$
$p(z)=\prod_{i\in\Omega}p(z_{i,1})\prod_{t=2}^{T}p(z_{i,t}|z_{i,t-1}),\;\;\mbox{with}$
$p(z_{i,t}=k|z_{i,t-1}=l)=\frac{1}{M}\exp(\alpha\mathbb{1}[k=l]+\beta\mathbb{1}[k\neq l])$ (7)
$\mbox{and}\;\;p(z_{i,1}=k)=\frac{1}{K},$ (8)
where $M$ is a normalizing factor. We introduce the more compact notation $g_{i,t}^{k}=g_{\phi}(f)_{i,t}^{k}$. Using an uninformative prior on $\theta$, the KL term is given by:
$KL[q(z,\theta|f)||p(z,\theta)]$ $=KL[q(z|f)||p(z)]$
$=\mathbb{E}_{q}[\log q(z|f)]-\mathbb{E}_{q}[\log p(z)].$ (9)
We then infer the two terms of eq.(9)
$\mathbb{E}_{q}$ $[\log q(z|f)]=$
$\sum_{i\in\Omega}\sum_{t=2}^{T}\sum_{k=1}^{K}\sum_{l=1}^{L}g_{i,t}^{k}g_{i,t-1}^{l}(\alpha\mathbb{1}[k=l]+\beta\mathbb{1}[k\neq l])+c,$ (10)
where $c=-|\Omega|(\log K+T\log M)$ is a constant. We set $\beta=0$, and knowing that $\alpha>0$, we have:
$\mathbb{E}_{q}[\log p(z)]=\alpha\sum_{i\in\Omega}\sum_{t=2}^{T}\sum_{k=1}^{K}g_{i,t}^{k}g_{i,t-1}^{k}+c.$ (11)
Since $2*g_{i,t}^{k}g_{i,t-1}^{k}=-(g_{i,t}^{k}-g_{i,t-1}^{k})^{2}+(g_{i,t}^{k})^{2}+(g_{i,t-1}^{k})^{2}$, after simple calculations, we get:
$KL$ $[q(z|f)||p(z)]=\frac{\alpha}{2}\sum_{i\in\Omega}\sum_{t=2}^{T}\sum_{k=1}^{K}(g_{i,t}^{k}-g_{i,t-1}^{k})^{2}$
$+\sum_{i\in\Omega}\sum_{t=1}^{T}\sum_{k=1}^{K}g_{i,t}^{k}*(\log(g_{i,t}^{k})-\frac{\alpha}{2}*g_{i,t}^{k})$
$-\frac{\alpha}{2}\sum_{i\in\Omega}\sum_{t=2}^{T-1}\sum_{k=1}^{K}(g_{i,t}^{k})^{2})+c.$ (12)
Since $\forall x\in[0,1],-\frac{\alpha}{2}\leq x(\log x-\frac{\alpha}{2}x)\leq 0$, we can infer:
As $\forall x,y\in[0,1]^{2},|x-y|\geq(x-y)^{2}$, we end up with the term of temporal consistency in the loss function we introduced in ST-MS [2]. Thus, we have provided a mathematical development to motivate it. More precisely, it writes:
$\mathcal{L}_{c}(f)=\frac{1}{2K|\Omega|}\sum_{i\in\Omega}\sum_{t=2}^{T}\sum_{k=1}^{K}\;|g_{\phi}(f)_{i,t}^{k}-g_{\phi}(f)_{i,t-1}^{k}|.$ (13)
The Eulerian temporal constraint expressed by eq.(13) does not mean that the site should be static between two successive frames, but it holds as long as the site lies within the successive projections of a given moving part so that it remains in the same mask. However, it does not make sense on occluded or disoccluded areas. This justifies the use of the L1 norm to deal with the latter configuration. We further prevent from enforcing the temporal consistency over occlusion areas by ignoring sites $i$ in the summation over $\mathcal{I}$ in eq.(13) that exhibit a large temporal flow difference. More precisely, for each input flow sequence, we set a threshold $\lambda$ so that a quantile $\eta$ of sites $i$ is discarded as follows:
$p(\|f(i,t+1)-f(i,t)\|_{1}\geq\lambda)\leq\eta.$ (14)
In practice, we take $\eta=1\%$ with an empirical distribution for $p$ in eq.(14). In doing so, we make an implicit assumption on the overall surface of the occlusion areas, but it seems reasonable for the datasets we deal with.
The loss function is finally defined by:
$\sum_{f\in\mathcal{D}}\mathcal{L}(f,\phi,\theta)=\sum_{f\in\mathcal{D}}\mathcal{L}_{r}(f,\phi,\theta)+\gamma\mathcal{L}_{c}(f,\phi).$ (15)
@@
We set $\gamma=1$. The training procedure alternatively updates $\theta$ and $\phi$ with $\epsilon$ as learning rate:
for $\text{for }f\in D:\;\;$ $\theta^{*}=\arg\min_{\theta}\mathcal{L}(f,\phi,\theta)(f)\;;\;$ (16)
$\phi_{t+1}=\phi_{t}-\epsilon\nabla_{\phi}\mathcal{L}(f,\phi,\theta^{*})$ (17)
@@ 3.3 Network architecture
The overall architecture of our unsupervised multiple motion segmentation framework is illustrated in Fig.1. It includes two main modules. The first one, taking the flow volume as input, is a 3D U-net [39]. The latent code augmented with embedding positions is directed to the transformer decoder. Then, by cross-attention, the output formed by the volume of segmentation masks is produced. The training of the overall architecture is based on the minimization of the loss function defined in eq.(15), while the motion model parameters of the segments are given by the B-spline estimation. Temporal consistency is provided by the loss function and the space-time motion models.
@@ 4.1 Implementation details
Following [7, 21, 29, 2, 49], we adopt the RAFT method [41] to compute the optical flow fields. More specifically, we use the RAFT version trained on the MPI Sintel dataset [3]. We downsample the computed flow fields to feed the network with 128$\times$224 vector fields. The output segmentation maps are upsampled to the initial frame size for evaluation. Thus, we can perform efficient training and inference stages. We typically take flow volumes of temporal length $T=9$ at training time. However, at test time, we can process flow volumes of larger temporal length, and even the flow volume of the whole sequence. To compute the spatio-temporal parametric motion models, $x$ and $y$ coordinates are normalized within $[-1,1]$, and a similar normalization is applied to the $t$ coordinate. The full 12-parameter quadratic motion model is used in all the experiments. We take $\nu=3$ for the frequency factor in the B-spline approximation. We select for each point $(x,y)$ the segment $\hat{k}$ with the highest prediction. In eq.(13), we set $\lambda$ as the $99^{th}$ quantile of the flow differences over the flow volume.
We use Adam optimizer with the following strategy on the learning rate $\epsilon$ to train the network (3D U-net and transformer), inspired from the warmup-decay strategy [47]. We linearly increase it from 00 to $41e-$ 4 for 20 epochs, then, we divide it by two every 40 epochs. The estimation of the motion model parameters through the B-spline approximation at training time is achieved with the Pytorch implementation of L-BGFS [22]. Indeed, the parametric models induced over time by the splines are a linear combination of quadratic parametric models (see eq.(2)). Thus, we can optimize eq.(3.2) to obtain the argmin in eq.(16), which is a convex problem.
Our LT-MS method is very efficient at test time (code made available upon publication). The computational time amounts on average to 210 fps on a GPU P100, that is, twice as fast as the ST-MS method [2]. It is due to the long flow sequence given as input to the network, which allows for parallelisation of some heavy computations. In addition, our LT-MS architecture remains lightweight, since it combines only three U-net layers and a transformer decoder on the downsampled feature space. Let us also stress that our LT-MS method does not involve any post-processing at all.
@@ 4.2 Data augmentation and network training
We introduced two types of data augmentation dedicated to optical flow. First, we add a 2D global flow to the input flow similarily as we did in the EM [29] and ST-MS [2] methods. However, for LT-MS, the 2D global flow is given by a full spline-based spatio-temporal motion model whose parameters are chosen at random. The same 2D global flow is added to the flow fields of a given input volume. It allows us to mimic diverse camera motions, enforcing that the motion segments are independent of the 2D added global flow, as justified in [29]. In addition, we corrupt a few flows out of the nine input ones. Thus, we simulate poorly estimated flow fields at some time instants, and the temporal consistency should overcome it.
Our motion segmentation method is fully unsupervised. We never use any manual segmentation annotation. We train our model using the FlyingThings3D (FT3D) dataset [27]. Our (demanding) objective is to demonstrate the high generalization power of our model. We train our model once for all on the optical flow fields computed on FT3D, and then, apply it on new unseen datasets. For hyperparameter setting, we select the stopping epoch from the loss function evaluated on the DAVIS2016 training set. We trained on a single Quadro RTX 6000 for 77h with a batch size of 5 on 32k flow fields computed on FT3D.
Ablation / Dataset DAVIS2016 FBMS59 SegTrackV2
Full Model LT-MS-K4 72.4 58.2 60.4
Unet3D only 71.3 55.5 57.3
No consistency term $\mathcal{L}_{c}$ 71.0 55.5 57.5
Polynomial space-time model 69.8 54.5 56.6
Table 1: Ablation study for three main components of our method LT-MS ($K=4$) on DAVIS2016, FBMS59 and SegTrackV2. Only one model component is modified at a time. Input sequence length is 120 (in practice, it comprises the whole video for the DAVIS2016 dataset). The metrics used is the Jaccard index $\mathcal{J}$. All evaluations are performed on the binary ground truth for the three datasets as explained in Section 5.5.
@@ 5 Experimental Results
Our LT-MS method is above all a long-term multiple motion segmentation method. However, in order to quantitatively assess its performance and in absence of OFS benchmarks, we need to resort to VOS benchmarks. We have carried out several types of experiments on four datasets: DAVIS2016111https://davischallenge.org/index.html [35], SegTrackV2222https://paperswithcode.com/dataset/segtrack-v2-1 [20], FBMS59 [32], and DAVIS2017-motion [47]. However, most VOS benchmarks are concerned with the segmentation of the primary moving object in the foreground, which means a binary ground truth. This led us first to test our LT-MS method in a two-mask ($K=2$) configuration, as discussed in Section 5.3, and reported in Table 2. Then, we have performed multiple motion segmentation evaluation, especially on DAVIS2017-motion, as commented in Section 5.4 and Table 3. Finally, we have assessed the long-term impact of LT-MS both on a quantitative (Table 4, Section 5.5) and qualitative (Section 5.6) basis.
Figure 3: Three groups of qualitative results regarding the ablation of the temporal-consistency loss term ($K=4$). They respectively correspond to the worm video of SegTrackV2, the dogs01 video of FBMS59, and the car-roundabout videos of DAVIS2016. For each group, the first row contains sample images with the segmentation ground-truth overlaid in yellow, the second row displays the input flows, the third and fourth rows show the predicted motion segmentations, respectively without and with the temporal consistency loss term. Clearly, this model component allows us to get far more consistent segments over time.
5.1 Datasets
DAVIS2016 comprises 50 videos (for a total of 3455 frames), depicting diverse moving objects. It is split in a training set of 30 videos and a validation set of 20 videos. Only the primary moving object is annotated in the ground truth. The criteria for evaluation on this dataset are the Jaccard score $\mathcal{J}$ (intersection-over-union), and the contour accuracy score $\mathcal{F}$, the higher, the better for both. FBMS59 includes 59 videos (720 annotated frames), and SegTrackV2 14 videos (1066 annotated frames). Both mostly involve one foreground moving object, but sometimes a couple of moving objects. For FBMS59, we use the 30 sequences of the validation set. Although annotations may comprise multiple objects, the VOS community exploits SegTrackV2 similarily as DAVIS2016, considering a binary ground-truth by grouping moving objects in the foreground. We follow this practice. However, we provide a multi-mask evaluation for FBMS59 as other works did.
@@ 5.2 Ablation study
We have conducted an ablation study to assess three main components of our long-term multiple-motion segmentation method LT-MS. To this end, we took four masks ($K=4$), and call LT-MS-K4 this version of our LT-MS method. These three components are mainly related to the temporal dimension of the problem. We have changed one component at a time as specified hereafter. We use the polynomial space-time quadratic motion model of ST-MS [2] instead of the space-time motion model based on B-splines over the input sequence. We omit the consistency term $\mathcal{L}_{c}$ in the loss function. We just take the convnet without the transformer decoder. All the ablation experiments were run on the three datasets, DAVIS2016, FBMS59, SegTrackV2. Results are collected in Table 1. We performed them for the input sequence length of 120 (in practice, the whole video for the DAVIS2016 dataset).
We can observe that the three ablations have almost the same impact on the performance. The three corresponding model components, i.e., the spline-based motion model, the temporal-consistency loss term, and the transformer decoder are thus all beneficial in similar proportions. They are able to handle the temporal dimension of the problem and the temporal motion evolution along the sequence in a compelling way. The contributions of these three components are even more significant for the FBMS59 and the SegTrackV2 datasets. In fact, the dynamic content of the majority of DAVIS2016 videos, along with performance scores calculated on a frame basis on these videos, do not allow to fully appreciate the contributions of the three model components. Yet, they can be highlighted by visualizing results on some videos of DAVIS2016, as shown in Fig.3.
Methods Training Input DAVIS2016 SegTrackV2
$\mathcal{J}$uparrow $\mathcal{J}$uparrow
Ours LT-MS-K2 Unsupervised Flow 70.7 58.9
EM [29] 69.3 55.5
MoSeg [49] 68.3 58.6
FTS [34] 55.8 47.8
TIS0 [12] 56.2 -
OCLR [47] (flow only) Additional Supervision 72.1 67.6
DyStab - Dyn [51] 62.4 40.0
GWM [7] Unsupervised RGB (Flow in loss) 79.5 78.3
RCF [21] 80.9 76.7
AMD [23] 57.8 57.0
MOD [10] 73.9 62.2
DivA(4$\dagger$ [1] RGB & Flow 72.4 64.6
IKE [13] 76.0 66.4
TISs [12] 62.6 -
CIS - No Post [50] 59.2 45.6
DS [52] Optimisation at test time 79.1 72.1
GO [36] Optimisation at test time & Additional Supervision 80.2 74.9
@@
5.3 Quantitative and comparative binary evaluation
We first report in Table 2 the quantitative evaluation on DAVIS2016 and SegTrackV2, performed on the binary ground truth (foreground moving object vs background) for the two datasets. This is the most frequent evaluation reported in the VOS literature. Therefore, we first apply our LT-MS method with two masks ($K=2$), denoted LT-MS-K2. Since our OFS does not establish the assignment of the two segments to respectively the background and the primary moving object, we rely on a simple heuristic designating the biggest segment as the background one.
Let us recall that our actual objective is the optical flow segmentation (OFS) task, even if we evaluate our method on VOS benchmarks in absence of OFS benchmarks. Since the two VOS benchmarks deal with the segmentation of the primary object moving in the foreground, it may occur discrepancies with OFS, which negatively impacts the scores for OFS methods. If wrong w.r.t. VOS ground truth, the segmentation of additional moving parts actually makes sense from the OFS standpoint, such as moving objects in the background, subparts of moving articulated bodies, or static foreground objects exhibiting motion parallax with a moving camera. Therefore, the comparative evaluation regarding OFS must be restricted to unsupervised methods taking only optical flow as input, namely, the top part of Table 2. Nevertheless, for a more comprehensive view, we also include results of VOS methods that leverage both RGB images and flow as input, in the bottom part of Table 2.
We evaluate LT-MS-K2 on a sequence basis, not on a frame basis as done by the other methods. This is more demanding, but we have done so to be consistent with the long-term intrinsic property of our method. We took a sequence length $T=120$ for DAVIS2016 and SegTrackV2. Table 2 also collects results obtained by other existing unsupervised methods also with two masks, when available on these benchmarks. We keep the categorization that we proposed in [29] regarding input and training. Nevertheless, we have added a category w.r.t. the network input for four recent methods, [7, 10, 21, 23], that only use RGB images as imput at test time, the optical flow being only involved in the loss function. OCLR leverages additional supervision, since it relies on human-annotated sprites to include realistic shapes in the computer-generated data used at training time. We consider the OCLR version taking only optical flow as input. The same holds for the DyStab method, the DyStab-Dyn version is considered. The post-processing added to the CIS method [50], based on Conditional Random Fields (CRF), is an heavy one, which has led most authors to retain only the version without post-processing for a fair comparison. We also include a separate category for two recent unsupervised methods, DS [52] and GO [36], where optimisation is performed at test time.